OpenStudy (anonymous):

find the product: (1/4x62-2/3)(1/4x-1/4) find the product: (1/4x62-2/3)(1/4x-1/4) @Mathematics

OpenStudy (anonymous):

(1/4x^2-2/3)(1/4x-1/4)

jhonyy9 (jhonyy9):

x are on numerators or on denominators ?

OpenStudy (anonymous):

let a=1/4 let b=3/4 (ax^2-b)(ax-a) a^2x^3-a^2x^2-abx+ab

OpenStudy (anonymous):

plug in the variables then your done.

OpenStudy (anonymous):

b=2/3 (correction)

jhonyy9 (jhonyy9):

a2x2(x-1) -ab(x-1)=(a2x2-ab)(x-1)=a(ax2-b)(x-1)

jhonyy9 (jhonyy9):

(1/4 x2 -2/3 )(1/4 x -1/4) =((3x2 -8)/12)((x-1)/4) =(3x2-8)/12 * 4/(x-1) = = (3x2 -8)/3 * 1/(x-1) = (3x2 -8)/(3(x-1))